i can not for the love of myself figure out how to solve the system equations by graphing.
3x+y=7
x+2y=4

Answer by pamalah@yahoo.com(10)

3x+y=7 becomes y=-3x+7
x+2y=4 becomes y=-1/2x+2
For equation 1 plot (0,7) and then plot second point by counting slope as down 3 right 1
For equation 2 plot (0,2) and then plot second point by counting slope as down 1 right 2